# A Comprehensive Evaluation on Quantization Techniques for Large Language Models

### v2

@@ 3.3 Granularity
Table 14: Evaluation of FP4 format quantization using low bit scaling factor and group size 16, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. E4M3, E2M1, and INT4 refer to the format of the scaling factor. W and A denote weight and activation, respectively. All of these follow NVFP4’s approach, employing a two-level scaling scheme.
We further reduce the bit-width of the scaling factor to 4 bits to evaluate its effectiveness under such a low-bit setting. As shown in Table 14, applying 4-bit scaling factors to both weights and activations leads to very poor performance. When applied only to weights, the performance remains moderate and is worse compared with NVFP4.
Evaluate Methods Wiki2‑PPL($\downarrow$)
Pre-quantization
transformation
Quantization error
mitigation
INT4
channel, FP32
MXFP4
g32, E8M0
NVFP4
g16, E4M3
scaling GPTQ 134.85 13.24 10.97
scaling low-rank 165.4 14.91 11.36
scaling GPTQ+low-rank 124.56 12.58 10.81
optimized rotation+scaling GPTQ 11.8 12.37 11.04
optimized rotation+scaling low-rank 14.58 15.43 12.57
optimized rotation+scaling GPTQ+low-rank 11.73 12.29 11.01
Table 15: Evaluation of INT4, MXFP4, and NVFP4 using both pre-quantization transformation and quantization error mitigation techniques, in terms of Perplexity on WikiText2 of LlaMA-3.2-1B. Channel means quantization granularity is per-channel; g16 and g32 denote per-group quantization with group sizes of 16 and 32, respectively. FP32, E8M0, and E4M3 refer to the format of the scaling factor.
Notably, rotation-based methods yield smaller improvements for MXFP4 and NVFP4 than for INT4 (per-channel). As argued in AMXFP4 [41], grouping activation tensors into small micro-scaled units mitigates outliers like rotation. For MXFP4 and NVFP4 formats, the group size is extremely small, so data rotation contributes little and may even lead to degradation.
4 Conclusions
In this paper, we have presented an experimental evaluation of Large language model quantization. We systematically decouple the commonly used quantization into two key steps: pre-quantization transformation and quantization error mitigation. After a comprehensive review of existing approaches, we choose scaling and rotation in pre-quantization transformation step, GPTQ, and low-rank compensation in the quantization error mitigation step to further conduct experimental evaluations on the same ground. We found that using the optimized rotation and scaling, together with two compensation methods, demonstrates the best performance. For quantization settings, asymmetric quantization and smaller group sizes yield better results. And using asymmetric quantization on activations shows more benefits on performance compared with weights. For FP4 quantization, we can adopt methods used in INT4 quantization. However, rotation-based pre-quantization transformation yields no observable more performance gains for MXFP4 and NVFP4 than scaling-based methods. And we find that the format of scaling factors demonstrates a large effect. For FP8 format scaling factor, FP8 (E4M3) with an additional per-tensor FP32 scaling factor is better than FP8 (E8M0). For small group sizes, using a 4-bit scaling factor for weights is also feasible, albeit with some loss in accuracy. To achieve optimal performance under FP4 quantization, further investigation is required to explore FP4’s full potential and limitations.

### v3

@@ III-C2 Exploration of FP4 scaling factor
IV Conclusions
In this paper, we have presented an experimental evaluation of Large language model quantization. We systematically decouple the commonly used quantization into two key steps: pre-quantization transformation and quantization error mitigation. After a comprehensive review of existing approaches, we choose scaling and rotation in the pre-quantization transformation step, GPTQ, and low-rank compensation in the quantization error mitigation step to further conduct experimental evaluations on the same ground. We found that using the optimized rotation and scaling, together with two compensation methods, demonstrates the best performance. For quantization settings, asymmetric quantization and smaller group sizes yield better results. And using asymmetric quantization on activations shows more benefits in performance compared with weights. For FP4 quantization, we can adopt methods used in INT4 quantization. However, rotation-based pre-quantization transformation yields no observable performance gains for MXFP4 and NVFP4 than scaling-based methods. And we find that the format of scaling factors demonstrates a large effect. For FP8 format scaling factor, FP8 (E4M3) with an additional per-tensor FP32 scaling factor is better than FP8 (E8M0). For small group sizes, using a 4-bit scaling factor for weights is also feasible, albeit with some loss in accuracy. To achieve optimal performance under FP4 quantization, further investigation is required to explore FP4’s full potential and limitations.